OpenStudy (anonymous):

If z=1/2 what is the value of z^-6z^4?

OpenStudy (anonymous):

\[z^{-6}z^4=z^{-6+4}=z^{-2}=\frac{1}{z^2}\]

OpenStudy (anonymous):

if you replace z by \[\frac{1}{2}\] you will get 4

OpenStudy (anonymous):

\[z^{-6}z^4=z^{-6+4}=z^{-2}\] If z=1/2, this becomes \[\left(\frac{1}{2}\right)^{-2}=(2^{-1})^{-2}=2^2=4\]